1242) y A x ln(x) + B x + F is the particular solution of the second-order linear DEQ: xy =2 where y=4 at the point (3,2)
Answer #1

1242 Solution Given that y= A x ln (x) + Bx + F is the particular solution of the second order and also given that 24=2 wher(from @ . W.K.T » 1 - - 9 Qn(3)+c >>C= 4-2 ln(3) =) B=(-2 > B: 4-286(3)-2 =) B=2-2 ln (3) -) B= -0.197 (substitute evalue)